OpenStudy (anonymous):

x(x+8)=0 zero product property

jimthompson5910 (jim_thompson5910):

The zero product property states that if AB = 0, then either A = 0 or B = 0 (or both are zero) So if x(x+8) = 0, then x = 0 or x+8 = 0. I'll let you take over.